# A parallel framework for graphical optimal transport

### v1

@@ Abstract
We study multi-marginal optimal transport (MOT) problems where the underlying cost has a graphical structure. These graphical multi-marginal optimal transport problems have found applications in several domains including traffic flow control and regression problems in the Wasserstein space. MOT problem can be approached through two aspects: a single big MOT problem, or coupled minor OT problems. In this paper, we focus on the latter approach and demonstrate it has efficiency gain from the parallelization. For tree-structured MOT problems, we introduce a novel parallelizable algorithm that significantly reduces computational complexity. Additionally, we adapt this algorithm for general graphs, employing the modified junction trees to enable parallel updates. Our contributions, validated through numerical experiments, offer new avenues for MOT applications and establish benchmarks in computational efficiency.
11footnotetext: This work was supported by the Swedish Research Council (VR) under grant 2014-5870, SJTU-KTH cooperation grant, the Knut and Alice Wallenberg foundation under grant KAW 2021.0274, and the NSF under grant 2206576.
@@ 1 Introduction
Optimal transport (OT) theory offers a fundamental mathematical framework for the comparison of probability distributions (Peyré et al. 2017). Central to OT is the quest for the most cost-efficient strategy to morph one distribution into another. With significant strides in the development of computationally efficient methods (Cuturi 2013, Zanetti and Gondzio 2023), OT has burgeoned into a plethora of applications. It has made substantial impacts in biology (Bunne et al. 2022, Demetci et al. 2020), generative modeling (Fan et al. 2021, 2022b, Korotin et al. 2022), image processing (Peyré et al. 2017, Fan et al. 2021, Zhou and Parno 2024), domain adaptation (Seguy et al. 2017, Perrot et al. 2016), Bayesian statistics (Fan et al. 2020, Srivastava et al. 2015), and control theory (Chen et al. 2016, 2021a, Singh et al. 2022, Ringh et al. 2021a), showcasing the versatility and the expansive scope of OT as a tool for analytical and computational advancement. Multi-marginal optimal transport (MOT) (Ba and Quellmalz 2022, Beier et al. 2023, Trillos et al. 2023, Yang and Zha 2023), a generalization of OT, involves optimizing over transports among more than two distributions, thereby extending the applicability of OT to a broader range of problems (Pass 2012, 2015, Nenna 2016). Its applications are vast and varied, encompassing areas such as signal processing (Elvander et al. 2020), estimation and control (Ringh et al. 2023, 2021a), and machine learning (Haasler et al. 2021c, Noble et al. 2023).
@@
The MOT problem frequently embodies a graphical structure where each node denotes a distribution, and the edges represent the transport plans. For example, the well-known Wasserstein barycenter problem finds a generalized average of a set of distributions (Agueh and Carlier 2011). This is done by finding a distribution that minimizes the sum of optimal transport distances to the given distributions, which can be represented by a star-shaped graph (Fan et al. 2022a). Originally, the Wasserstein barycenter problem was formulated as a sum of coupled bi-marginal OT problems (Carlier and Ekeland 2010, Rabin et al. 2012). However, it can equivalently be posed as a MOT problem, where the graphical structure is instead invoked through the underlying cost. This equivalence between the MOT formulation and the coupled OT formulation holds for any structured problems, where the underlying graph is a tree, i.e., an acyclic graph (Haasler et al. 2021b). Although the coupled OT methodology often presents a more intuitive formulation, its computational development has not been extensively pursued, and existing numerical solvers for general graphical structures rely predominantly on the MOT formulation (Elvander et al. 2020, Haasler et al. 2021b, c, Fan et al. 2022a). To close this gap in the literature, we develop numerical methods for graph-structured OT problems that rely on the coupled OT formulation.
@@
More precisely, following the powerful standard approach of adding an entropic term to the OT objective (Cuturi 2013), we first develop a novel scheme for a sum of OT problems that are coupled according to a tree-structure. We also prove complexity results for solving the problem to a desired accuracy. It turns out that this algorithm can be easily parallelized, thus significantly improving the runtime of previous algorithms that are based on the equivalent MOT problem. We also extend our algorithm to the setting of general graph-structures that may contain cycles. In this setting, the algorithm is based on factorizing the underlying graph as a modified junction tree and solving an OT problem that is structured according to this modified junction tree. Thus, in this case the problem is posed as a sum of several coupled MOT problems of small size, where each MOT problem corresponds to a clique, i.e., a collection of nodes, in the modified junction tree.
To summarize, our contributions are the following.
@@
(i) For MOT problems associated with tree structures, we reformulate them as coupled bi-marginal OT problems and introduce a novel algorithm for solving them with local entropy regularization in Sec. 4.1. In our approach, we represent the tree as a bipartite graph and categorize the optimization variables into two groups. This partitioning facilitates the parallelization of our algorithm. This advancement bridges the existing gap in the literature by providing a dedicated algorithmic solution for coupled small OT problems.
@@
(ii) In Sec. 4.2, we conduct a thorough complexity analysis for our algorithm addressing coupled bi-marginal OT problems within tree structures. Contrasting with standard MOT problems that iterate over nodes sequentially, our parallelizable algorithm requires only $\widetilde{\mathcal{O}}\left({|E|^{2}}/{\delta^{2}}\right)$ iterations. Here, $|E|$ denotes the total number of edges, and $\delta$ specifies the desired accuracy. This represents a significant efficiency gain, reducing the number of iterations by a factor of $|E|$ compared to the standard MOT problem approach analyzed in Fan et al. (2022a).
@@
(iii) In Sec. 5, we introduce the notion of a modified junction tree, allowing us to extend the scope of our algorithm from tree structures to general graphs. This extension involves formulating the problem as a sum of coupled small MOT problems and constructing a bipartite graph to enable a parallel algorithm. To validate the efficacy and robustness of our algorithm, we have conducted several numerical experiments, the results of which affirm the algorithm’s applicability to graph structures.
(iv) In Sec. 3, we present a concise framework for MOT problems with global entropy regularization. Extending our previous work Fan et al. (2022a), this framework allows for marginal constraints on the joint distributions. Furthermore, it incorporates the capacity for cost functions to depend on multiple marginals, a feature previously implied but now explicitly defined and integrated into our algorithm.
@@ 2.1.1 Optimal transport
The OT problem seeks an optimal transport plan that moves mass from a source distribution to a target one with minimum cost. In this work, we focus on the discrete OT problem (see, e.g., Villani (2003) for the continuous formulations). Thus, consider two non-negative vectors $\mu_{1}\in{\mathbb{R}}_{+}^{d}$ and $\mu_{2}\in{\mathbb{R}}_{+}^{d}$ with mass $1$, i.e., $\sum_{x_{1}}\mu_{1}(x_{1})=\sum_{x_{2}}\mu_{2}(x_{2})=1$, that describe discrete probability distributions. For simplicity of notation, we assume the marginals have equal dimension $d$ throughout, though the extension to the setting with unequal dimensions is straightforward. Here $\mu_{1}(x)$ and $\mu_{2}(x)$ denote the amount of mass in the source distribution and target distribution, respectively, at location $x$. We also define a cost matrix $C\in{\mathbb{R}}^{d\times d}$, where $C(x_{1},x_{2})$ describes the transport cost of moving a unit mass from point $x_{1}$ to $x_{2}$. Similarly, we describe the transport by a non-negative matrix $B\in{\mathbb{R}}_{+}^{d\times d}$, where the element $B(x_{1},x_{2})$ describes the amount of mass transported from point $x_{1}$ to $x_{2}$. This defines the optimal transport problem
$\min_{B\in{\mathbb{R}}_{+}^{d\times d}}{}$ $\operatorname{trace}(C^{\top}B)\qquad\text{ subject to }$ $B{\bf 1}=\mu_{1},\quad B^{\top}{\bf 1}=\mu_{2},$ (1)
@@ 2.1.2 Multi-marginal optimal transport
The MOT problem extends the OT problem 1 to finding a transport plan between multiple distributions $\mu_{1},\dots,\mu_{J}$, where $J>2$. Let $\mu_{j}\in{\mathbb{R}}^{d}$ for $j=1,\dots,J$. Then the multi-marginal transport cost and transport plan are described by $J$-mode tensors ${\bf C}\in{\mathbb{R}}^{d\times\dots\times d}$ and ${\bf M}\in{\mathbb{R}}_{+}^{d\times\dots\times d}$. Here, ${\bf C}(x_{1},x_{2},\ldots,x_{J})$ and ${\bf M}(x_{1},x_{2},\ldots,x_{J})$ denote the cost and amount of transport associated with the locations $x_{1},x_{2},\ldots,x_{J}$. Thus, analogously to 1, the MOT problem is defined as
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j},\text{ for }j\in\Gamma,$ (2)
@@
$P_{j}({\mathbf{B}})=\sum_{x_{1},\dots,x_{j-1},x_{j+1},x_{J}}{\mathbf{B}}(x_{1},\dots,x_{j-1},x_{j},x_{j+1},\dots,x_{J}).$ (3)
In the original MOT formulation (Gangbo and Świech 1998, Pass 2015), constraints are given on all marginal distributions, that is the index set in the constraints is $\Gamma=\{1,2,\dots,J\}$. However, as we will see in the next section, in many applications not all marginal distributions are given, thus in this work $\Gamma$ is typically a strict subset of $\{1,2,\dots,J\}$. Finally, we note that the standard bi-marginal OT problem 1 is a special case of the MOT problem 2 with $J=2$ and $\Gamma=\{1,2\}$.
@@ 2.1.3 Iterative scaling algorithm
We note that MOT is a standard linear program. However, as it has $d^{J}$ variables, solving it directly scales exponentially with the number of marginals $J$. The computational burden of solving 2 can be partly alleviated by regularizing the objective with an entropy term. This approach has first been proposed for the bi-marginal problem 1 in Cuturi (2013), Chen et al. (2021b), and has later been adapted to multi-marginal problems (Benamou et al. 2015, Haasler et al. 2021b, c). The entropy regularized MOT problem is the strictly convex optimization problem
$\min_{{\mathbf{B}}\in{\mathbb{R}}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}|{\bf M})\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j},\text{ for }j\in\Gamma,$ (4)
@@
${\mathcal{H}}({\mathbf{B}}|{\bf M})=\sum_{x_{1},\dots,x_{J}}{\mathbf{B}}(x_{1},\dots,x_{J})\left(\log\left({\mathbf{B}}(x_{1},\dots,x_{J})\right)-\log\left({\bf M}(x_{1},\dots,x_{J})\right)-1\right).$ (5)
with ${\bf M}(x_{1},\dots,x_{J})=\prod_{j\in\Gamma}\mu_{j}(x_{j})$. The entropy regularized problem 4 can be solved by an iterative scaling algorithm, presented in Algorithm 1. This method is also known as Sinkhorn iterations, especially in the bi-marginal setting 1.
@@
In the following, we briefly sketch how to derive Algorithm 1 as a block coordinate ascent method in the dual of 4, for details we refer the reader to Elvander et al. (2020), Haasler et al. (2021b, c). Namely, utilizing Lagrangian duality theory, one can show that the optimal solution to 4 is of the form
${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M},$ (6)
where $\odot$ denotes element-wise multiplication and the tensors are given by
${\bf K}=\exp(-{\bf C}/\epsilon)\qquad{\bf U}=u_{1}\otimes u_{2}\otimes\dots\otimes u_{J},$ (7)
where the vectors $u_{j}\in\mathbb{R}^{d}$ are given by
$u_{j}=\begin{cases}\exp\left(\frac{\lambda_{j}}{\epsilon}\right),&\text{ if }j\in\Gamma\\ {\bf 1},&\text{ otherwise,}\end{cases}$ (8)
and $\lambda_{j}\in\mathbb{R}^{d}$ is the dual variable corresponding to the constraint $P_{j}({\mathbf{B}})=\mu_{j}$ on the $j$-th marginal. The corresponding dual problem of 4 is
$\max_{\{\lambda_{j},j\in\Gamma\}}-\epsilon\langle{\bf K},{\bf U}\rangle{+\sum_{j\in\Gamma}\lambda_{j}^{\top}\mu_{j},}$ (9)
where ${\bf U}$ depends on the dual variables as in 7-8. Algorithm 1 is then derived as a block coordinates ascent in this dual and is therefore guaranteed to converge globally.
@@ 2.2 Graph-structured optimal transport
${\bf C}({\bf x})=\sum_{\alpha\in F}C_{\alpha}({\bf x}_{\alpha}),$ (10)
where $F$ is a set of subsets of the tensors indices, i.e., $\alpha\subset\{1,\dots,J\}$ for each element $\alpha\in F$. If we associate the index set of the transport tensor with a set of vertices, $V=\{1,\dots,J\}$, then the set $F$ can be understood as the set of factors in a factor graph (Haasler et al. 2021c, Fan et al. 2022a). In particular, all subsets $\alpha\in F$ that contain two elements can be illustrated as edges in a graph. Moreover, we allow for constraints that act on projections onto several marginals of the transport tensor ${\mathbf{B}}$. For a set $\alpha\subset V$, we define the projection $P_{\alpha}({\mathbf{B}})=\sum_{x_{j},\ j\in V\backslash\alpha}{\mathbf{B}}(x_{1},\dots,x_{J}).$ Let $F_{\Gamma}\subset F$ denote the set of factors which are constrained, and for $\alpha\in F_{\Gamma}$, let $\boldsymbol{\mu}_{\alpha}\in{\mathbb{R}}_{+}^{d^{|\alpha|}}$ be the given tensor, where the number of modes equals the cardinality $|\alpha|$ of $\alpha$. This let us define the graph-structured MOT problem as
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{\alpha}({\mathbf{B}})=\boldsymbol{\mu}_{\alpha},\text{ for }\alpha\in F_{\Gamma},$ (11)
where the cost tensor decouples as in 10, and $F_{\Gamma}\subset F$. Note that the graph structured MOT problem 11 is a special case of the standard MOT problem 2 in terms of the cost function. On the other hand, problem 11 allows for a more general class of constraints than the original problem 2.
Remark 1.
@@
The most widely studied graph-structured MOT problem is the Wasserstein barycenter problem (Agueh and Carlier 2011). However, in a recent line of research, graph-structured MOT problems have been utilized in a wide range of applications. For instance, in Haasler et al. (2021b, a) different control and estimation problems for large systems of indistinguishable agents are modeled as graph-structured MOT problems. Even different classes of agents can be modeled with this framework, e.g., in network flow problems (Haasler et al. 2024) and mean field games (Ringh et al. 2021b, 2023). In the following we briefly review a few more exampled of optimal transport problems that fall into the category of graph-structured MOT. In particular, we note that to the best of our knowledge, the Euler problem and Wasserstein splines have not yet been classified as such. Throughout this paper, we illustrate constrained marginals as gray nodes, and white nodes are estimated in the respective problems.

### v2

@@ Abstract
We study multi-marginal optimal transport (MOT) problems where the underlying cost has a graphical structure. These graphical multi-marginal optimal transport problems have found applications in several domains including traffic flow control, barycenter and regression problems in the Wasserstein space, and Hidden Markov model inference problems. The MOT problem can be approached through two formulations: a single big MOT problem, or coupled minor OT problems. In this paper, we focus on the latter approach and demonstrate its efficiency gain from parallelization. For tree-structured MOT problems, we introduce a novel parallelizable algorithm that significantly reduces computational complexity. Additionally, we adapt this algorithm for general graphs, employing the modified junction trees to enable parallel updates. Our contributions, validated through numerical experiments, offer new avenues for MOT applications and establish benchmarks in computational efficiency.
keywords:
Optimal transport, Probabilistic graphical model, Parallel algorithms, Optimization, Complexity analysis
{MSCcodes}
49M29; 65K05; 90C06
@@ 1 Introduction
Optimal transport (OT) theory offers a fundamental mathematical framework for the comparison of probability distributions [peyre2017computational]. Central to OT is the quest for the most cost-efficient strategy to morph one distribution into another. With significant strides in the development of computationally efficient methods [Cut13, zanetti2023interior], OT has burgeoned into a plethora of applications. It has made substantial impacts in control theory [CheGeoPav14e, CheGeoPav21a, SinHaaZha20, ringh2021efficient], generative modeling [fan2021neural, fan2021variational, korotin2022neural], image processing [peyre2017computational, fan2021neural, zhou2022efficient], domain adaptation [seguy2017large, perrot2016mapping], Bayesian statistics [fan2020scalable, srivastava2015wasp], and biology [bunne2022proximal, demetci2020gromov], showcasing the versatility and the expansive scope of OT as a tool for analytical and computational advancement. Multi-marginal optimal transport (MOT) [ba2022accelerating, beier2023unbalanced, trillos2023multimarginal, yang2023estimating], a generalization of OT, involves optimizing over transports among more than two distributions, thereby extending the applicability of OT to a broader range of problems [Pas12, Pas15, Nen16]. Its applications are vast and varied, encompassing areas such as estimation and control [ringh2023mean, ringh2021efficient], signal processing [ElvHaaJakKar20], and machine learning [HaaSinZha20, noble2023tree].
@@
The MOT problem frequently embodies a graphical structure where each node denotes a distribution, and the edges represent the transport plans. For example, the well-known Wasserstein barycenter problem finds a generalized average of a set of distributions [AguCar11]. This is done by finding a distribution that minimizes the sum of optimal transport distances to the given distributions, which can be represented by a star-shaped graph [fan2021complexity]. Originally, the Wasserstein barycenter problem was formulated as a sum of coupled bi-marginal OT problems [carlier2010matching, rabin2012wasserstein]. However, it can equivalently be posed as a MOT problem, where the graphical structure is instead invoked through the underlying cost. This equivalence between the MOT formulation and the coupled OT formulation holds for any structured problems, where the underlying graph is a tree, i.e., an acyclic graph [HaaRinChe20]. Although the coupled OT methodology often presents a more intuitive formulation, its computational development has not been extensively pursued, and existing numerical solvers for general graphical structures rely predominantly on the MOT formulation [ElvHaaJakKar20, HaaRinChe20, HaaSinZha20, fan2021complexity]. To close this gap in the literature, we develop numerical methods for graph-structured OT problems that rely on the coupled OT formulation.
@@
More precisely, following the powerful standard approach of adding an entropic term to the OT objective [Cut13], we first develop a novel scheme for a sum of OT problems that are coupled according to a tree-structure. We also prove complexity results for solving the problem to a desired accuracy. It turns out that this algorithm can be easily parallelized, thus significantly improving the runtime of previous algorithms that are based on the equivalent MOT problem. We also extend our algorithm to the setting of general graph-structures that may contain cycles. In this setting, the algorithm is based on factorizing the underlying graph as a modified junction tree and solving an OT problem that is structured according to this modified junction tree. Thus, in this case the problem is posed as a sum of several coupled MOT problems of small size, where each MOT problem corresponds to a clique, i.e., a collection of nodes, in the modified junction tree. To summarize, our contributions are the following.
@@
(i) For MOT problems associated with tree structures, we reformulate them as coupled bi-marginal OT problems and introduce a novel algorithm for solving them with local entropy regularization in Section 3.1. In our approach, we represent the tree as a bipartite graph and categorize the optimization variables into two groups. This partitioning facilitates the parallelization of our algorithm. This advancement bridges the existing gap in the literature by providing a dedicated algorithmic solution for coupled small OT problems.
@@
(ii) In Section 3.2, we conduct a thorough complexity analysis for our algorithm addressing coupled bi-marginal OT problems within tree structures. Contrasting with standard MOT problems that iterate over nodes sequentially, our parallelizable algorithm requires only $\widetilde{\mathcal{O}}\left({|E|^{2}}/{\delta^{2}}\right)$ iterations. Here, $|E|$ denotes the total number of edges, and $\delta$ specifies the desired accuracy. This represents a significant efficiency gain, reducing the number of iterations by a factor of $|E|$ compared to the standard MOT problem approach analyzed in [fan2021complexity].
@@
(iii) In Section 4, we introduce the notion of a modified junction tree, allowing us to extend the scope of our algorithm from tree structures to general graphs. This extension involves formulating the problem as a sum of coupled small MOT problems and constructing a bipartite graph to enable a parallel algorithm. To validate the efficacy and robustness of our algorithm, we have conducted several numerical experiments, the results of which affirm the algorithm’s applicability to graph structures.
@@ 2.1.1 Optimal transport
The OT problem seeks an optimal transport plan that moves mass from a source distribution to a target one with minimum cost. In this work, we focus on the discrete OT problem (see, e.g., [Vil03] for the continuous formulations). Thus, consider two non-negative vectors $\mu_{1}\in{\mathbb{R}}_{+}^{d}$ and $\mu_{2}\in{\mathbb{R}}_{+}^{d}$ with mass $1$, i.e., $\sum_{x_{1}}\mu_{1}(x_{1})=\sum_{x_{2}}\mu_{2}(x_{2})=1$, that describe discrete probability distributions. For simplicity of notation, we assume the marginals have equal dimension $d$ throughout, though the extension to the setting with unequal dimensions is straightforward. Here $\mu_{1}(x)$ and $\mu_{2}(x)$ denote the amount of mass in the source distribution and target distribution, respectively, at location $x$. We also define a cost matrix $C\in{\mathbb{R}}^{d\times d}$, where $C(x_{1},x_{2})$ describes the transport cost of moving a unit mass from point $x_{1}$ to $x_{2}$. Similarly, we describe the transport by a non-negative matrix $B\in{\mathbb{R}}_{+}^{d\times d}$, where the element $B(x_{1},x_{2})$ describes the amount of mass transported from point $x_{1}$ to $x_{2}$. This defines the optimal transport problem
(1) $\min_{B\in{\mathbb{R}}_{+}^{d\times d}}$ $\operatorname{trace}(C^{\top}B)\qquad\text{ subject to }$ $B{\bf 1}=\mu_{1},\quad B^{\top}{\bf 1}=\mu_{2},$
@@ 2.1.2 Multi-marginal optimal transport
The MOT problem extends the OT problem Eq. 1 to finding a transport plan between multiple distributions $\mu_{1},\dots,\mu_{J}$, where $J>2$. Let $\mu_{j}\in{\mathbb{R}}^{d}$ for $j=1,\dots,J$. Then the multi-marginal transport cost and transport plan are described by $J$-mode tensors ${\bf C}\in{\mathbb{R}}^{d\times\dots\times d}$ and ${\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}$. Here, ${\bf C}(x_{1},x_{2},\ldots,x_{J})$ and ${\mathbf{B}}(x_{1},x_{2},\ldots,x_{J})$ denote the cost and amount of transport associated with the locations $x_{1},x_{2},\ldots,x_{J}$. Thus, analogously to Eq. 1, the MOT problem is defined as
(2) $\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j},\text{ for }j\in\Gamma,$
@@
(3) $P_{j}({\mathbf{B}})=\sum_{x_{1},\dots,x_{j-1},x_{j+1},x_{J}}{\mathbf{B}}(x_{1},\dots,x_{j-1},x_{j},x_{j+1},\dots,x_{J}).$
In the original MOT formulation [GanSwi98, Pas15], constraints are given on all marginal distributions, that is the index set in the constraints is $\Gamma=\{1,2,\dots,J\}$. However, as we will see in the next section, in many applications not all marginal distributions are given, thus in this work $\Gamma$ is typically a strict subset of $\{1,2,\dots,J\}$. Finally, we note that the standard bi-marginal OT problem Eq. 1 is a special case of the MOT problem Eq. 2 with $J=2$ and $\Gamma=\{1,2\}$.
@@ 2.1.3 Iterative scaling algorithm
We note that MOT is a standard linear program. However, as it has $d^{J}$ variables, solving it directly scales exponentially with the number of marginals $J$. The computational burden of solving Eq. 2 can be partly alleviated by regularizing the objective with an entropy term. This approach has first been proposed for the bi-marginal problem Eq. 1 in [Cut13, CheGeoPav21b], and has later been adapted to multi-marginal problems [BenCarCut15, HaaRinChe20, HaaSinZha20]. The entropy regularized MOT problem is the strictly convex optimization problem
(4) $\min_{{\mathbf{B}}\in{\mathbb{R}}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}|{\bf M})\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j},\text{ for }j\in\Gamma,$
@@
(5) ${\mathcal{H}}({\mathbf{B}}|{\bf M})=\!\!\!\sum_{x_{1},\dots,x_{J}}\!\!\!{\mathbf{B}}(x_{1},\dots,x_{J})\left(\log\left({\mathbf{B}}(x_{1},\dots,x_{J})\right)-\log\left({\bf M}(x_{1},\dots,x_{J})\right)-1\right)\!\!$
with ${\bf M}(x_{1},\dots,x_{J})=\prod_{j\in\Gamma}\mu_{j}(x_{j})$. The entropy regularized problem Eq. 4 can be solved by an iterative scaling algorithm, presented in Algorithm 1. This method is also known as Sinkhorn iterations, especially in the bi-marginal setting Eq. 1.
@@
In the following, we briefly sketch how to derive Algorithm 1 as a block coordinate ascent method in the dual of Eq. 4, for details we refer the reader to [ElvHaaJakKar20, HaaRinChe20, HaaSinZha20]. Namely, utilizing Lagrangian duality theory, one can show that the optimal solution to Eq. 4 is of the form ${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M}$, where $\odot$ denotes element-wise multiplication and the tensors are given by
(6) ${\bf K}=\exp(-{\bf C}/\epsilon,)\quad{\bf U}=u_{1}\otimes u_{2}\otimes\dots\otimes u_{J},\quad u_{j}=\begin{cases}\exp\left(\frac{\lambda_{j}}{\epsilon}\right),&\text{ if }j\in\Gamma\\ {\bf 1},&\text{ otherwise,}\end{cases}$
where $\lambda_{j}\in\mathbb{R}^{d}$ is the dual variable corresponding to the constraint $P_{j}({\mathbf{B}})=\mu_{j}$ on the $j$-th marginal. The corresponding dual problem of Eq. 4 is
(7) $\max_{\{\lambda_{j},j\in\Gamma\}}-\epsilon\langle{\bf K},{\bf U}\rangle{+\sum_{j\in\Gamma}\lambda_{j}^{\top}\mu_{j},}$
where $\bf U$} depends on the dual variables as in Eq. 6. Algorithm 1 is then derived as a block coordinates ascent in this dual and is therefore guaranteed to converge globally.
@@ 2.2 Graph-structured optimal transport
(8) ${\bf C}({\bf x})=\sum_{\alpha\in F}C_{\alpha}({\bf x}_{\alpha}),$
where $F$ is a set of subsets of the tensors indices, i.e., $\alpha\subset\{1,\dots,J\}$ for each element $\alpha\in F$. If we associate the index set of the transport tensor with a set of vertices, $V=\{1,\dots,J\}$, then the set $F$ can be understood as the set of factors in a factor graph [HaaSinZha20, fan2021complexity]. In particular, all subsets $\alpha\in F$ that contain two elements can be illustrated as edges in a graph. Moreover, we allow for constraints that act on projections onto several marginals of the transport tensor $\mathbf{B}$}. For a set $\alpha\subset V$, we define the projection $P_{\alpha}({\mathbf{B}})=\sum_{x_{j},\ j\in V\backslash\alpha}{\mathbf{B}}(x_{1},\dots,x_{J}).$ Let $F_{\Gamma}\subset F$ denote the set of factors which are constrained, and for $\alpha\in F_{\Gamma}$, let $\bm{\mu}_{\alpha}\in{\mathbb{R}}_{+}^{d^{|\alpha|}}$ be the given tensor, where the number of modes equals the cardinality $|\alpha|$ of $\alpha$. This let us define the graph-structured MOT problem as
(9) $\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{\alpha}({\mathbf{B}})=\bm{\mu}_{\alpha},\text{ for }\alpha\in F_{\Gamma},$
where the cost tensor decouples as in Eq. 8, and $F_{\Gamma}\subset F$. Note that the graph structured MOT problem Eq. 9 is a special case of the standard MOT problem Eq. 2 in terms of the cost function. On the other hand, problem Eq. 9 allows for a more general class of constraints than the original problem Eq. 2.
Remark 2.1.
@@
The most widely studied graph-structured MOT problem is the Wasserstein barycenter problem [AguCar11]. However, in a recent line of research, graph-structured MOT problems have been utilized in a wide range of applications. For instance, in [HaaRinChe20, haasler2021control] different control and estimation problems for large systems of indistinguishable agents are modeled as graph-structured MOT problems. Even different classes of agents can be modeled with this framework, e.g., in network flow problems [haasler2023scalable] and mean field games [ringh2021graph, ringh2023mean]. In the following we briefly review a few more exampled of optimal transport problems that fall into the category of graph-structured MOT. In particular, we note that to the best of our knowledge, the Euler problem and Wasserstein splines have not yet been classified as such. Throughout this paper, we illustrate constrained marginals as gray nodes, and white nodes are estimated in the respective problems.